# Zak-OTFS over CP-OFDM

### v1

@@ I Introduction
For example, numerical simulations in Section VII for the TDL-C channel (Urban-macro, normal delay profile and a maximum Doppler shift of $1.25$ KHz) for a packet having $1$ ms duration and bandwidth $720$ KHz, at a total received signal power to noise power ratio (TSNR) of $14$ dB (including data and pilots), reveal that the effective SE of Zak-OTFS over CP-OFDM (after considering the pilot and CP overheads) is $40$ percent more than that achieved with CP-OFDM. Therefore, even with existing CP-OFDM based modems it is possible to achieve a significant improvement in the effective SE by simply implementing Zak-OTFS over CP-OFDM using the proposed architecture. The effective SE of unconstrained Zak-OTFS is about $6\%$ higher than that of Zak-OTFS over CP-OFDM. For higher channel Doppler spreads, the gains are even higher. As an example, for a maximum Doppler shift of $6.48$ KHz and a line-of-sight (LOS) TDL-D channel (Urban-macro, short delay profile [27]), the effective SE of CP-OFDM, Zak-OTFS over CP-OFDM and unconstrained Zak-OTFS at a TSNR of $14$ dB are respectively, $0.56$, $1.33$ and $1.92$ bits/sec/Hz. Note that the effective SE achieved with unconstrained Zak-OTFS is roughly $44$ percent more than that achieved with Zak-OTFS over CP-OFDM. A maximum Doppler shift of $1.25$ KHz is possible for a carrier frequency of $15$ GHz and a mobile speed of $90$ Km/hr (vehicular speed). A higher speed of $1000$ Km/hr (e.g., aircraft-to-ground communication) at a carrier frequency of $7$ GHz results in a maximum Doppler shift of $6.48$ KHz.
@@ II Preliminaries: Doubly-spread channel, CP-OFDM and Zak-OTFS modulation
$h(\tau,\nu)$ $=$ $\sum\limits_{i=1}^{P}h_{i}\,\delta(\tau-\nu_{i})\,\delta(\nu-\nu_{i}),$ (2)
@@ II-B Zak-OTFS modulator and demodulator
$x_{dd}[k,l]$ $=$ $x[k,l]\,e^{j2\pi\lfloor\frac{k}{M}\rfloor\frac{l}{N}},$ (19)
@@ II-C Discrete Zak Transform (DZT) based implementation of Zak-OTFS
Result 2 (see Appendix A in [12] and also [24])
@@ VII Numerical results
We consider communication in the Frequency Range 3 (FR3) band with carrier frequency $f_{c}=15,7$ GHz, and possible sub-carrier spacings (SCS) $\Delta f=15,30,60$ KHz. We simulate three scenarios as listed in Table I. Scenarios-I and II model non-line-of-sight (Non-LOS) communication with a user in a vehicle (car) and train respectively. Scenario-III models LOS aircraft-to-ground (A2G) communication with an aircraft speed of $1000$ km/hr (IMT-2030 considers mobile speeds up to $1000$ Km/hr).
@@ II $15$ GHz TDL-C $216$ 33 KHz Normal
Another interesting observation in Fig. 12 is that for a TSNR of $14$ dB, unconstrained Zak-OTFS achieves an SE which is roughly $44$ percent more than that achieved by Zak-OTFS over CP-OFDM with $N=1$. This is due to the freedom in choosing the filters and the Zak-OTFS period parameters. This shows the potential of achieving significant improvements in SE by not constraining Zak-OTFS numerology and filters.

### v2

@@ I Introduction
For example, numerical simulations in Section VII for the TDL-C channel (Urban-macro, normal delay profile and a maximum Doppler shift of $1.25$ KHz) for a packet having $1$ ms duration and bandwidth $720$ KHz, at a total received signal power to noise power ratio (TSNR) of $14$ dB (including data and pilots), reveal that the effective SE of Zak-OTFS over CP-OFDM (after considering the pilot and CP overheads) is $40$ percent more than that achieved with CP-OFDM. Therefore, even with existing CP-OFDM based modems it is possible to achieve a significant improvement in the effective SE by simply implementing Zak-OTFS over CP-OFDM using the proposed architecture. The effective SE of unconstrained Zak-OTFS is about $6\%$ higher than that of Zak-OTFS over CP-OFDM. For higher channel Doppler spreads, the gains are even higher. As an example, for a maximum Doppler shift of $6.48$ KHz and a line-of-sight (LOS) TDL-D channel (Urban-macro, short delay profile [27]), the effective SE of CP-OFDM, Zak-OTFS over CP-OFDM and unconstrained Zak-OTFS at a TSNR of $14$ dB are respectively, $0.56$, $1.33$ and $1.92$ bits/sec/Hz. Note that the effective SE achieved with unconstrained Zak-OTFS is roughly $30$ percent more than that achieved with Zak-OTFS over CP-OFDM. A maximum Doppler shift of $1.25$ KHz is possible for a carrier frequency of $15$ GHz and a mobile speed of $90$ Km/hr (vehicular speed). A higher speed of $1000$ Km/hr (e.g., aircraft-to-ground communication) at a carrier frequency of $7$ GHz results in a maximum Doppler shift of $6.48$ KHz.
We have also simulated a no-mobility scenario (zero Doppler shift) with a long delay profile (maximum delay spread of roughly $10\mu s$). For the proposed Zak-OTFS over CP-OFDM, the highest SE is achieved when $M=1$ (i.e., minimum delay period) which is CP-OFDM itself. However, unconstrained Zak-OTFS achieves significantly better performance than CP-OFDM. This is because, inter-symbol interference (ISI) degrades the effective SE achieved by CP-OFDM as the CP duration (with 3GPP 5G NR numerology) is limited to $4.7\mu s$ which is not sufficient for large delay profile scenarios. For example, at a total received signal to noise ratio of $14$ dB, Zak-OTFS (unconstrained) achieves almost double the effective SE achieved with CP-OFDM. This scenario highlights the importance of Zak-OTFS not just for high mobility scenarios but also for long delay scenarios.
@@ II Preliminaries: Doubly-spread channel, CP-OFDM and Zak-OTFS modulation
$h(\tau,\nu)$ $=$ $\sum\limits_{i=1}^{P}h_{i}\,\delta(\tau-\tau_{i})\,\delta(\nu-\nu_{i}),$ (2)
@@ II-B Zak-OTFS modulator and demodulator
$x_{dd}[k,l]$ $=$ $x[k\,\text{mod}\,M,l\,\text{mod}\,N]\,e^{j2\pi\lfloor\frac{k}{M}\rfloor\frac{l}{N}},$ (19)
@@ II-C Discrete Zak Transform (DZT) based implementation of Zak-OTFS
Result 2 (see Chapter $8$ (Appendix A) in [12] and also [24])
@@ VII Numerical results
We consider communication in Frequency Range 3 (FR3) band with carrier frequency $f_{c}=15,7$ GHz, and also Frequency Range 1 (FR1). Possible standardized sub-carrier spacings (SCS) for CP-OFDM are $\Delta f=15,30,60$ KHz. We simulate four scenarios as listed in Table I. Scenarios-I and II model non-line-of-sight (NLOS) communication with a user in a vehicle (car) and train respectively. Scenario-III models LOS aircraft-to-ground (A2G) communication with an aircraft speed of $1000$ km/hr (IMT-2030 considers mobile speeds up to $1000$ Km/hr). Scenario IV models NLOS for a long delay profile with no mobility $\nu_{max}=0$ Hz, $f_{c}=3.5$ GHz.
@@ II $15$ GHz TDL-C $216$ 33 KHz Normal
Another interesting observation in Fig. 12 is that for a TSNR of $14$ dB, unconstrained Zak-OTFS ($\nu_{p}=24$ KHz) achieves an SE which is roughly $30$ percent more than that achieved by Zak-OTFS over CP-OFDM with $N=1$. This is due to the freedom in choosing the filters and the Zak-OTFS period parameters. This shows the potential of achieving significant improvements in SE by not constraining Zak-OTFS numerology and filters.
Figure 13: Effective spectral efficiency vs. total received signal power to noise ratio (TSNR) (dB), for Scenario-IV (no mobility, long delay profile) (see Table I).
In Fig. 13 we plot the effective SE for Scenario-IV (TDL-C, long delay profile in Table 7.7.3-2 of [27], no mobility). The maximum delay spread is $9.93\,\mu s$. Since, there is no mobility, for CP-OFDM it suffices to choose a $15$ KHz SCS. For unconstrained Zak-OTFS it suffices to choose the minimum possible Doppler period which satisfies the conditions in (57) and (58). Since $\nu_{p}=N/T$ and $T=1$ ms, the minimum possible Doppler period which satisfies (57) and (58) is $\nu_{p}=1/T=1$ KHz which corresponds to $N=1$ and $M=720$. Since the unconstrained Zak-OTFS pilot overhead fraction is $19N/720$, $N=1$ also corresponds to minimum pilot overhead.
In Fig. 13 it is observed that the effective SE performance of CP-OFDM ($15$ KHz SCS) is significantly inferior to that of unconstrained Zak-OTFS (with $\nu_{p}=1$ KHz). This is due to the degradation in CP-OFDM resulting from inter-symbol interference (ISI) as the CP duration of $4.7\mu s$ is less than the delay spread of $9.93\mu s$. Note that CP-OFDM is nothing but Zak-OTFS over CP-OFDM with minimum delay period ($M=1$). For this no mobility scenario, we observed that the highest SE was achieved with $M=1,N=48$ and so we have not plotted the curves for other $(M,N)$ configurations.